### v2

@@ B.1 Hyperparameter Analysis
B.2 Hyperparameters settings
All experiments in this paper are conducted on a GPU of Nvidia GeForce RTX 3090 and a CPU of AMD EPYC 7742. Our implementation of SAC, MEow, TD3, QSM, DACER, QVPO, and DIPO follows their official codes: https://github.com/toshikwa/soft-actor-critic.pytorch, https://github.com/ChienFeng-hub/meow, https://github.com/sfujim/TD3, https://github.com/Alescontrela/score_matching_rl, https://github.com/happy-yan/DACER-Diffusion-with-Online-RL, https://github.com/wadx2019/qvpo, and https://github.com/BellmanTimeHut/DIPO. The shared hyperparameters of all algorithms are listed in Table $2$When comparing with other diffusion-based algorithms, MaxEntDP uses 3-layer MLPs as the actor and critic networks following the default settings of these algorithms. In other experiments, 2-layer MLPs are used as they can already attain good performance..
Table 2: The shared hyperparameters of all algorithms.
Hyperparameter MaxEntDP SAC MEow TD3 QSM DACER QVPO DIPO
Batch size 256 256 256 256 256 256 256 256
Discount factor $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
No. of hidden layers 2 2 2 2 3 3 3 3
No. of hidden nodes 256 256 256 256 256 256 256 256
Actor learning rate 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4
Critic learning rate 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4
Activation mish relu relu relu mish mish mish mish
Replay buffer size 1e6 1e6 1e6 1e6 1e6 1e6 1e6 1e6
Diffusion steps 20 N/A N/A N/A 20 20 20 20
Action candidate number 10 N/A N/A N/A N/A N/A 32 N/A
B.3 Training time
The training time for all algorithms is presented in Table 3. Leveraging the computation efficiency of JAX [48] and the parallel processing capabilities of GPU, MaxEntDP demonstrates high training efficiency compared to competing methods, only behind TD3 and QSM. This highlights its advantage for real-world applications that require high computation efficiency.
Algorithm (2-layer MLP) MaxEntDP (jax) SAC MEow TD3
Training time (h) 3.9 4.6 11 1.7
Training time (h) 5.5 1.9 5.9 22.6 55.6

### v3

@@ D.1 Hyperparameter Analysis
D.2 Multimodal Policy Learning on the Challenging AntMaze Benchmarks
Figure 8: Trajecories generated by MaxEntDP and SAC after 1M environment interactions in Antmaze benchmarks.
Figure 9: State coverage of MaxEntDP and SAC after 1M environment interactions in Antmaze benchmarks. MaxEntDP can explore different behavior modes at the same time and show broader state coverage than SAC, exhibiting efficient exploration of the high-dimensional state-action space.
We adopt the AntMaze benchmarks proposed in DDiffPG (Li et al., 2024) to test the multi-modal policy learning ability of MaxEntDP on the challenging high-dimensional RL tasks. In this environment, the agent is a quadruped robot trying to reach the specified goals. Instead of the sparse reward employed in DDiffPG, we use a dense reward, a penalty for the distance from the closest goal, to guide policy learning. We demonstrate the trajectories generated by MaxEntDP and SAC after 1M environment interactions in Figure 8. MaxEntDP can learn diverse behavior modes even in the challenging high-dimensional tasks, while SAC fails to learn different solutions. In addition, we visualize state coverage through the training process for MaxEntDP and SAC, showing the results in Figure 9. We can see that MaxEntDP can explore multiple behavior modes at the same time, while SAC focuses only on a simple mode. This reveals the importance of using the expressive diffusion policy for efficient exploration and multimodal policy learning.
D.3 Comparative Evaluation on the DeepMind Control Suite
Figure 10: Learning curves on DeepMind Control suite. The solid lines are the means, and the shaded regions represent the standard deviations over five runs.
We test MaxEntDP on 3 high-dimensional tasks on the DeepMind Control Suite benchmarks. The performance comparison with SAC, MEow, and TD3 is displayed in Figure 10. MaxEntDP outperforms other baseline algorithms on these challenging high-dimensional RL tasks.
D.4 Testing the Accuracy of the Proposed Numerical Integration Technique on Probability Approximation
In theory, the numerical integration will be accurate when the diffusion step $T$ and the number of samples $N$ for probability approximation become large enough, according to the Law of Large Numbers. To exhibit the accuracy of different $T$ and $N$, we conduct experiments on a simple 2-D toy example where the target distribution $p(x)$ is a mixture of four Gaussian distributions with equal weights. Therefore, we set $Q(x)=\log p(x)$ and utilize the QNE method proposed in our paper to train a diffusion model. We display the approximation results of different $T$ and $N$ in Figure 11. The setting $T=20,N=50$ used in the paper can provide an effective probability approximation for the diffusion policy. When the samples are less ($T=20,N=20$), although there is a non-negligible error to the ground truth, the numerical integration method can still assign higher values for the region with higher probability. In this case, the estimated log probability can be considered as a kind of intrinsic curiosity reward to promote the exploration of the action region with low policy probability.
Figure 11: The probability approximation using numerical integration method on a 2-D toy example with different diffusion steps $T$ and sample numbers $N$. The target distribution is a mixture of four Gaussian distributions, whose means are (-0.5, -0.5), (-0.5, 0.5), (0.5, 0.5) and (0.5, -0.5). The standard deviations and weights of the four components are the same, which are 0.1 and 0.25, respectively. The setting in the paper ($T=20,N=50$) can provide an effective approximation for the true log probability. When fewer samples ($T=20,N=10$) are used, despite some estimation errors, our method can still assign higher values to high-probability regions.